# Finite Measures of Maximal Entropy for an Open Set of Partially Hyperbolic Diffeomorphisms

### v1

@@ 1.5. Structure of the paper and the sketch of the proof of the Main Theorem
    Utilizing the assumption $h^{u}(f)>h^{s}(f)$, we established that every ergodic measure with substantial entropy possesses a negative central Lyapunov exponent that remains uniformly away from zero. That is, if $h_{\mu}(f)>h^{u}(f)$ then
    $\lambda_{c}(f,\mu)\leq h_{\mu}(f)-h^{u}(f),$
    where $\lambda_{c}(f,\mu)$ denotes the central Lyapunov exponent with respect to $\mu$.
@@
    We prove that every point in $\Lambda_{\rho}(f)$ possesses an invariant Pesin manifold of uniform size. Particularly, for two points in $\Lambda_{\rho}(f)$ that are sufficiently close, their invariant manifolds have transversal intersections.
@@
    Utilizing the classical Pliss Lemma, we establish that for every ergodic measure with sufficiently large entropy, the set $\Lambda_{\rho}(f)$ possesses a positive measure with respect to this measure.
Items (1), (2) and (3) imply that there is only a finite number of equivalence classes in the set of hyperbolic measures containing measures of maximal entropy (we are referencing the equivalence relation described in Definition 2.5). By Theorem 2.8, each of these equivalence classes possesses at most one measure of maximal entropy. Therefore, the number of maximal entropy ergodic measures for $f$ is finite.
@@
    We observe that the property $h^{u}(f)>h^{s}(f)$ is an open property in $C^{1+}$ diffeomorphisms on $M$, that is, for any $h^{u}(f)>a>b>h^{s}(f)$, there exists a $C^{1}$ neighborhood of $f$ such that every $C^{1+}$ diffeomorphism has unstable entropy greater than $a$ and stable entropy less than $b$.
@@
    For every diffeomorphism $g$ close enough to $f$ we establish an injective correspondence between the ergodic measure of maximal entropy of $g$ into the ergodic measure of maximal entropy of $f$.
The final item demonstrates that any $g$ sufficiently close to $f$ has, at most, the same number of ergodic measures of maximal entropy as $f$. Establishing an injective map between the ergodic measures of maximal entropy of $g$ and those of $f$ was the most intricate part of our work. This proof relies on two notable facts:
  • (a)
    Consider a sequence of $C^{1+}$ diffeomorphisms $(g_{n})_{n\in\mathbb{N}}$ converging to $f$ in the $C^{1}$ topology. If $\mu_{n}$ represents ergodic m.m.e’s for $g_{n}$ and $\mu_{n}$ converges to $\nu$, we prove that $\nu$ is a measure of maximal entropy for $f$.
  • (b)
    For $\mu_{n}$ and $\nu$ as in Item (a) we show that $\nu(\Lambda_{\rho}(f))>0$
Item (a) implies that $\nu$ forms a convex combination of ergodic measures of maximal entropy for $f$. Item (b) suggests that at least one measure within this convex combination corresponds to an ergodic measure of maximal entropy for $f$ gives a positive measure to certain subset of $\Lambda_{\rho}(f)$. This specific measure is the one we associate with $\mu_{n}$.
@@
In Section 2 we recall the main previous results that we will use. In section 3 we explore the consequences of the hypotheses $h^{u}(f)>h^{s}(f)$. The section 4 is dedicated to proving that the set of points with Pesin manifolds of uniform size has a uniform positive size for measures with large entropy. In sections 5 and 6 we prove the results of our work.
@@ 2.1. Hyperbolic sets and homoclinic class
Let $f:M\to M$ be a diffeomorphism on a compact manifold $M$, and let $\Lambda\subset M$ be a compact and invariant set. We say that $\Lambda$ is a hyperbolic set for $f$ if there exists a splitting of the tangent bundle $T\Lambda$ into two bundles $E^{s}$ and $E^{u}$ such that for all $x\in\Lambda$, $Df_{x}E^{s}_{x}=E^{s}_{f(x)}$ and $Df_{x}E^{u}_{x}=E^{s}_{f(x)}$ and there exist a constant $C>0$ and $10<\lambda<$ 1 such that for all $x\in\Lambda$ it holds
@@
$\left\|Df^{-n}_{x}v\right\|\leq C\lambda^{n}\left\|v\right\|\mbox{for all }v\in E^{u}_{x}\mbox{ and }n\geq 1$
@@
The sets $W^{s/u}(x)$ are called the stable and unstable manifolds of $x$ respectively, and for $\varepsilon$ small, $W^{s/u}_{\varepsilon}(x):=W^{s/u}(x)\cap B_{\varepsilon}(x)$ are called the local stable and unstable manifolds of $x$. When a periodic point is hyperbolic we say that its orbit is a hyperbolic periodic orbit. We denote the set of all hyperbolic periodic orbits of $f$ by $\operatorname{Per}_{h}(f)$. The transverse intersection of $C^{r}$-submanifolds $U,V\subset M$ is
$U\pitchfork V=\{x\in U\cap V:T_{x}U\oplus T_{x}V=T_{x}M\}$
@@
A hyperbolic set $\Lambda$ is called locally maximal if it possesses an open neighborhood $V$ satisfying $\Lambda=\bigcap_{n\in{\mathbb{Z}}}f^{n}(V)$. A fundamental property of hyperbolic sets is the existence of a $\delta>0$ such that, for any $x,y\in\Lambda$ with $d(x,y)<\delta$, the intersection $W_{\varepsilon}^{s}(x)\pitchfork W_{\varepsilon}^{u}(x)$ is a unique point $y\in M$. When this point $y$ belongs to $\Lambda$, we refer to $\Lambda$ as possessing a local product structure. It is known that a hyperbolic set $\Lambda$ is locally maximal if and only if it presents this property of local product structure [30].
@@ 2.2. Hyperbolic measures and homoclinic classes
We will recall some of the results presented in [2]. Let $x$ be a point of $M$. For a nonzero vector $v\in T_{x}M$, the Lyapunov exponent of $f$ at $v$ in $x$ is defined by
$\chi(x,v)=\lim_{{n\to+\infty}}\frac{1}{n}\log\left\|Df^{n}(v)\right\|$
@@
Let ${\mathcal{M}}_{f}^{e}$ o set of the ergodic measures of $f$. If $\mu\in{\mathcal{M}}_{f}^{e}$, the Oseledets theorem establishes that for almost every $x\in M$ with respect to $\mu$, the Lyapunov exponents of every nonzero vector $v\in T_{x}M$ exist [22].
@@
    $\chi(x,v)=\lim_{{n\to\pm\infty}}\frac{1}{n}\log\left\|Df^{n}_{v}\right\|=\chi_{i}.$
The set of points for which the Oseledest Theorem holds for a certain ergodic measure is called a Lyapunov regular set, denoted by ${{\mathcal{R}}}$.
@@
When $\mu\in{\mathcal{M}}_{f}^{h}$ we define for $\mu$-almost every $x\in M$ the splitting $T_{x}M=E^{+}_{x}(f)\oplus$ $E^{-}_{x}(f)$ where $E^{+}_{x},E^{-}_{x}$ are linear spaces such that $\lim_{n\rightarrow\infty}\frac{1}{n}\log\left\|Df_{x}^{n}v\right\|<0$ on $E^{+}(x)\backslash\{0\}$, and $\lim_{n\rightarrow\infty}\frac{1}{n}\log\left\|Df_{x}^{-n}v\right\|>0$ on $E^{-}(x)\backslash\{0\}$.
@@
$W^{-}(x)=\left\{y\in M:\limsup_{n\to+\infty}\frac{1}{n}\log d\left(f^{n}(x),f^{n}(y)\right)=0\right\}.$
@@
$W^{+}(x)=\left\{y\in M:\limsup_{n\rightarrow+\infty}\frac{1}{n}\log d\left(f^{-n}(x),f^{-n}(y)\right)=0\right\}.$
It is known that given $\mu\in{\mathcal{M}}_{f}^{h}$, for $\mu$ almost every $x\in M$ the stable and unstable Pesin sets are immersed submanifolds with the dimension ${\operatorname{dim}}(E^{-})$ and ${\operatorname{dim}}(E^{+})$ respectively and therefore are called stable and unstable Pesin manifolds [24]. Given a hyperbolic measure $\mu$ of $f$ we define the s-index of $\mu$ as s-index$(\mu):=\operatorname{Dim}(E^{+})$. In the remaining of the section, we consider only ergodic measures with the same s-index.
@@
$HC_{T}(\mu):=\overline{\bigcup\{\operatorname{supp}(\nu):\nu\in{\mathcal{M}}_{f}^{h},\mu\sim\nu\}}$
The relation $\left({\mathcal{M}}_{f}^{h},\sim\right)$ is an equivalence relation [2, Proposition 2.11].
@@
[2, Corollary 2.14.] For any $f\in{{\rm Diff}}^{r}(M)$, $r>1$, and any $\mu\in{\mathcal{M}}_{f}^{h}$, the set of measures supported by periodic orbits is dense in the set of hyperbolic ergodic measures homoclinically related to $\mu$, endowed with the weak-normal-$\ast$ topology. In particular, there exists ${{\mathcal{O}}}\in{\mathcal{M}}_{f}^{h}$ such that ${{\mathcal{O}}}\sim\mu$ and
$HC_{T}(\mu)=\overline{\bigcup\left\{\operatorname{supp}\nu:\nu\in{\mathcal{M}}_{f}^{h},\nu\sim{{\mathcal{O}}}\right\}}=HC(\mathcal{O}).$
@@
Let $f:M\rightarrow M$ be a $C^{r}$ diffeomorphism $r>1,\phi:M\rightarrow{\mathbb{R}}$ be a Hölder potential, and $p$ a hyperbolic periodic point with orbit ${{\mathcal{O}}}$. Then there is at most one, hyperbolic, equilibrium state for $\phi$ which is homoclinically related to ${{\mathcal{O}}}$, and its support coincides with ${{\rm HC}}({{\mathcal{O}}})$.
@@
The proof is a consequence of [23], which is a generalization of [29], and Theorem [2, Theorem 3.1]. ∎
@@
It is important to observe that the above proposition does not imply that for each $\mathcal{O}\in\operatorname{Per}(f)$, $HC(\mathcal{O})$ has at most one hyperbolic measure of maximal entropy. It can exist two orbits $\mathcal{O}_{1}$ and $\mathcal{O}_{2}\in\operatorname{Per}(f)$ that are not homoclinically related to each other, each one related to a hyperbolic measure for maximal entropy, such that $HC(\mathcal{O}_{1})\subset HC(\mathcal{O}_{2})$. Consequently, $HC(\mathcal{O}_{2})$ could have two ergodic measures of maximal entropy.
2.3. Partial hyperbolic systems
@@ 2.3.1. Existence of equilibrium states
$P_{metric}(f,\phi):{\mathcal{M}}_{f}\to\mathbb{R},\,\,\mu\mapsto P_{\mu}(f,\phi)$
@@
2.3.2. Hyperbolic measures
Consider a partially hyperbolic diffeomorphism $f$ with a central dimension one, denoted as $E^{c}$, and stable and unstable subspaces $E^{s}$ and $E^{u}$, respectively, and $E^{u}$ with dimensions greater than zero. Let $\mu$ be an ergodic measure associated to $f$. The measure $\mu$ is a hyperbolic measure of saddle type for $f$ if and only if for $\mu$-almost every point $x$ in $M$, the Lyapunov exponent in the central direction is different to zero, that is
$\lambda_{c}(\mu,f):=\lim_{n\to\infty}\frac{1}{n}\log|Df^{n}_{E^{c}}(x)|=\int\log|Df^{n}_{E^{c}}|d\mu\neq 0.$
Therefore, in the context of a partially hyperbolic system with a one-dimensional central component, determining whether a given ergodic measure is hyperbolic is achieved by verifying if the Lyapunov exponent in the central direction is non-zero
@@ 2.4. Unstable entropy
In this section, we recall the notions of unstable $h^{u}(f)$ and stable $h^{s}(f)$ topological entropy for a partial hyperbolic diffeomorphism $f$ introduced in [33, 10, 32].
@@ 2.4.1. Unstable metric entropy
Consider a compact metric space $M$ with the Borel $\sigma$-algebra ${{\mathcal{B}}}$ and $\mu$ represent a Borel probability measure defined on this space. Let ${{\mathcal{P}}}=\{P_{i}\}_{i\in I}$ be a partition of $M$ into measurable subsets, i.e., $M=\bigcup_{P_{i}\in\mathcal{P}}P_{i}$ such that $P_{i}\in{{\mathcal{B}}}$, $\mu(P_{i}\cap P_{j})=0$, and $\mu\left(\bigcup_{P_{i}\in{{\mathcal{P}}}}P_{i}\right)=\mu(M)=1$. We denote by $\sigma({{\mathcal{P}}})$ the smallest $\sigma$-algebra containing all $P_{i}\in{{\mathcal{P}}}$.
@@
We say that a partition ${{\mathcal{P}}}$ is measurable (or countably generated) with respect to $\mu$ if there exists a measurable family $\{A_{i}\}_{i\in\mathbb{N}}$ and a measurable set $F$ of full measure such that if $B\in{{\mathcal{P}}}$, then there exists a sequence $\{B_{i}\}$, where $B_{i}\in\{A_{i},A_{i}^{c}\}$ such that $B\cap F=\bigcap_{i}(B_{i}\cap F)$.
Consider $\pi:M\rightarrow M/P$ as the canonical projection, mapping each point in $M$ to the corresponding element in the quotient space formed by the atoms of the partition. This naturally gives rise to a measuring structure on $M/P$ via the measure $\widetilde{\mu}=\pi^{*}\mu$ defining it on the associated pushed $\sigma$-algebra. Additionally, for any point $x$ in $M$, $P(x)$ represents the element of the partition to which $x$ belongs.
@@
[26, Rokhlin’s Disintegration] Let $P$ be a measurable partition of a compact metric space $M$, and $\mu$ a Borel probability measure. Then there exists a unique disintegration by conditional probability measures for $\mu$.
@@
$h_{\mu}(P,f)=\lim_{n\to\infty}\frac{1}{n}H_{\mu}(P_{n})=\lim_{n\to\infty}\left(H_{\mu}(P_{n})-H_{\mu}(P_{n-1})\right),$
where $P_{n}=P\vee f^{-1}(P)\vee\ldots\vee f^{-n+1}(P)$ is a refinement of $P$ by the action of $f$.
Now let ${{\mathcal{F}}}$ be an $f$-invariant foliation, that is, $f{{\mathcal{F}}}(x))={{\mathcal{F}}}(f(x))$, for every $x\in M$. We say that ${{\mathcal{F}}}$ is uniformly expanding if there exists $\lambda>1$ such that $\forall x\in M$, $\|Df^{-1}_{|T_{x}(F(x))}\|<\lambda^{-1}$. A measurable partition $\xi$ is called increasing and subordinated to ${{\mathcal{F}}}$ if it satisfies the following properties:
@@
Finding a measurable and expanding partition that aligns with an invariant lamination is typically a complex challenge. However, when dealing with a uniformly expanding foliation that remains invariant under a diffeomorphism, such a partition is always available [33]. Additionally, for expanding flotations ${{\mathcal{F}}}$ it holds
@@
It can be proved that the above definition is independent of $\xi$.
@@
$h^{s}_{\mu}(f):=h_{\mu}(f,{{\mathcal{F}}}^{s}),$
where ${{\mathcal{F}}}^{s}$ denotes the unstable foliation given by the unstable manifolds.

### v3

@@ 1.5. Sketch of the proof of the Main Theorem
    Utilizing the assumption $h^{u}(f)>h^{s}(f)$, we established that every ergodic measure with substantial entropy possesses a negative central Lyapunov exponent that remains uniformly away from zero. That is, if $h_{\mu}(f)>h^{u}(f)$ then $\lambda_{c}(f,\mu)\leq h_{\mu}(f)-h^{u}(f),$ where $\lambda_{c}(f,\mu)$ denotes the central Lyapunov exponent with respect to $\mu$.
@@
    For every $\rho$ big enough, the set is not empty and every $x\in\Lambda_{\rho}(f)$ has invariant Pesin manifolds of uniform size. In particular, for two points in $\Lambda_{\rho}(f)$ sufficiently close, their invariant manifolds have transversal intersections.
@@
    Utilizing the classical Pliss Lemma, we prove that for certain values of $\rho$ (depending on $\lambda_{c}$), $\Lambda_{\rho}(f)$ has large measure for all invariant measure with large enough entropy, the set $\Lambda_{\rho}(f)$ has uniform positive measure with respect to this measure depending on how big is the entropy.
Items (1), (2) and (3) imply that there is only a finite number of homoclinic classes in the set of hyperbolic measures containing measures of maximal entropy (we are referencing the equivalence relation described in Definition 2.5). By Theorem 2.8, each of these equivalence classes possesses at most one measure of maximal entropy. Therefore, the number of maximal entropy ergodic measures for $f$ is finite.
@@
    Thanks to Lemma 3.4 the property $h^{u}(f)>h^{s}(f)$ is an open property in $C^{1+}$ diffeomorphisms on $M$, that is, for any $h^{u}(f)>a>b>h^{s}(f)$, there exists a $C^{1}$ neighborhood of $f$ such that every $C^{1+}$ diffeomorphism has unstable entropy greater than $a$ and stable entropy less than $b$.
@@
    To conclude the proof, we argue by contradiction. Assuming that there is a sequence of $C^{1+}$ diffeomorphisms $g_{n}$ having a number of ergodic measures of maximal entropy greater than the number of ergodic measures of maximal entropy of $f$ which converges to $f$. For each $g_{n}$ we take an ergodic measure $\nu_{n}$ of maximal entropy. There is no loss of generality assuming that $\nu_{n}$ converges to an $f$-invariant measure $\nu$.
  • (7)
    Using the consequences from our hypotheses $h^{u}(f)>h^{s}(f)$, we prove that $\nu$ is a measure of maximal entropy of $f$.
  • (8)
    We further get that $\nu$ has an ergodic component that we associate with $\nu_{n}$. In this way we establish an injective correspondence between the ergodic measure of maximal entropy of $g_{n}$ into the ergodic measure of maximal entropy of $f$. Thus the number of ergodic maximal entropy measures of $g_{n}$ is bounded by the number of ergodic maximal entropy measures of $f$, contradicting the hypotheses.
@@
In Section 2 we recall the main previous results that we will use. In section 3 we explore the consequences of the hypotheses $h^{u}(f)>h^{s}(f)$. The section 4 is dedicated to proving that the set of points with Pesin manifolds of uniform size has a uniform positive size for measures with large entropy. In sections 5 to 7 we prove the results of our work.
@@ 2.1. Hyperbolic sets and homoclinic class
Let $f:M\to M$ be a diffeomorphism on a compact manifold $M$, and let $\Lambda\subset M$ be a compact and invariant set. We say that $\Lambda$ is a hyperbolic set for $f$ if there exists a splitting of the tangent bundle $T\Lambda$ into two bundles $E^{s}$ and $E^{u}$ such that for all $x\in\Lambda$, $Df_{x}E^{s}_{x}=E^{s}_{f(x)}$ and $Df_{x}E^{u}_{x}=$ $E^{u}_{f(x)}$ and there exist a constant $C>0$ and $10<\lambda<$ 1 such that for all $x\in\Lambda$ it holds
@@
$\left\|Df^{-n}_{x}v\right\|\leq C\lambda^{n}\left\|v\right\|\mbox{for all }v\in E^{u}_{x}\mbox{ and }n\geq 1.$
@@
The sets $W^{s/u}(x)$ are called the stable and unstable manifolds of $x$ respectively, and for $\varepsilon$ small, we denote by $W^{s/u}_{\varepsilon}$ the connected component of $W^{s/u}\cap B_{\varepsilon}(x)$ that contains $x$. When a periodic point is hyperbolic we say that its orbit is a hyperbolic periodic orbit. We denote the set of all hyperbolic periodic orbits of $f$ by $\operatorname{Per}_{h}(f)$. The transverse intersection of $C^{r}$-submanifolds $U,V\subset M$ is
$U\pitchfork V=\{x\in U\cap V:T_{x}U\oplus T_{x}V=T_{x}M\}.$
@@
A hyperbolic set $\Lambda$ is called locally maximal if it possesses an open neighborhood $V$ satisfying $\Lambda=\bigcap_{n\in{\mathbb{Z}}}f^{n}(V)$. A fundamental property of hyperbolic sets is the existence of a $\delta>0$ such that, for any $x,y\in\Lambda$ with $d(x,y)<\delta$, the intersection $W_{\varepsilon}^{s}(x)\pitchfork W_{\varepsilon}^{u}(y)$ is a unique point $\ z\in M$. When this point $z$ belongs to $\Lambda$, we refer to $\Lambda$ as possessing a local product structure. It is known that a hyperbolic set $\Lambda$ is locally maximal if and only if it presents this property of local product structure [46].
@@ 2.2. Hyperbolic measures and homoclinic classes
We will recall some of the results presented in [4]. Let $x$ be a point of $M$. For a nonzero vector $v\in T_{x}M$, the Lyapunov exponent of $f$ $x$ on direction $v$ is defined by
$\chi(x,v):=\lim_{{n\to+\infty}}\frac{1}{n}\log\left\|Df^{n}(v)\right\|$
@@
Let ${\mathcal{M}}_{f}^{e}$ be the set of ergodic measures of $f$. If $\mu\in{\mathcal{M}}_{f}^{e}$, the Oseledets theorem establishes that for almost every $x\in M$ with respect to $\mu$, the Lyapunov exponent of every nonzero vector $v\in T_{x}M$ exist [35].
@@
    $\chi(x,v)=\lim_{{n\to\pm\infty}}\frac{1}{n}\log\left\|Df^{n}_{v}\right\|=\chi_{i}(f,\mu).$
The set of points for which the Oseledets Theorem holds for a certain ergodic measure is called a Lyapunov regular set, denoted by ${{\mathcal{R}}}$.
@@
When $\mu\in{\mathcal{M}}_{f}^{h}$ we define, for $\mu$-almost every $x\in M$, the splitting $T_{x}M=E^{+}_{x}(f)\oplus E^{-}_{x}(f)$ where $E^{+}_{x},\,E^{-}_{x}$ are linear spaces such that $\lim_{n\rightarrow\infty}\frac{1}{n}\log\left\|Df_{x}^{n}v\right\|<0$ on $E^{+}(x)\backslash\{0\}$, and $\lim_{n\rightarrow\infty}\frac{1}{n}\log\left\|Df_{x}^{-n}v\right\|>0$ on $E^{-}(x)\backslash\{0\}$.
@@
$W^{-}(x)=\left\{y\in M:\limsup_{n\to+\infty}\frac{1}{n}\log d\left(f^{n}(x),f^{n}(y)\right)<0\right\}.$
@@
$W^{+}(x)=\left\{y\in M:\limsup_{n\rightarrow+\infty}\frac{1}{n}\log d\left(f^{-n}(x),f^{-n}(y)\right)<0\right\}.$
It is known that given $\mu\in{\mathcal{M}}_{f}^{h}$, for $\mu$ almost every $x\in M$ the stable and unstable Pesin sets are immersed submanifolds with the dimension ${\operatorname{dim}}(E^{-})$ and ${\operatorname{dim}}(E^{+})$ respectively and therefore are called stable and unstable Pesin manifolds [39]. When $\{x,f(x)\cdots f^{p-1}(x)\}\in\operatorname{Per}_{h}(f)$ we have that $W^{s}(f^{i}(x))=W^{-}(f^{i}(x))$. Given a hyperbolic measure $\mu$ of $f$ we define the $\operatorname{s-index}$ of $\mu$ as $\operatorname{s-index}(\mu):=\operatorname{dim}(E^{-})$. In the remaining of the section, we consider only ergodic measures with the same s-index.
@@
$HC_{T}(\mu):=\overline{\bigcup_{\nu\in HC(\mu)}\operatorname{supp}(\nu)}.$
The relation $\left({\mathcal{M}}_{f}^{h},\sim\right)$ is an equivalence relation [4, Proposition 2.11].
@@
[4, Corollary 2.14.] For any $f\in{{\rm Diff}}^{r}(M)$, $r>1$, and any $\mu\in{\mathcal{M}}_{f}^{h}$, the set of measures supported by periodic orbits is dense in the set of hyperbolic ergodic measures homoclinically related to $\mu$, endowed with the weak-$\ast$ topology. In particular, there exists ${{\mathcal{O}}}\in\operatorname{Per}^{h}(f)$ such that ${{\mathcal{O}}}\sim\mu$ and
$HC_{T}(\mu)=HC(\mathcal{O}).$
@@
Let $f:M\rightarrow M$ be a $C^{r}$ diffeomorphism for some $r>1,\phi:M\rightarrow{\mathbb{R}}$ be a Hölder continuous potential, and $p$ a hyperbolic periodic point with orbit ${{\mathcal{O}}}$. Then there is at most one hyperbolic equilibrium state for $\phi$ which is homoclinically related to ${{\mathcal{O}}}$, and its support coincides with ${{\rm HC}}({{\mathcal{O}}})$.
@@
The proof is a consequence of [36], which is a generalization of [45], and Theorem [4, Theorem 3.1]. ∎
@@
It is important to observe that the above proposition does not imply that for each $\mathcal{O}\in\operatorname{Per}(f)$, $HC(\mathcal{O})$ has at most one hyperbolic measure of maximal entropy. There can exist two orbits $\mathcal{O}_{1}$ and $\mathcal{O}_{2}\in\operatorname{Per}(f)$ that are not homoclinically related to each other, each one related to a hyperbolic measure for maximal entropy, such that $HC(\mathcal{O}_{1})\subset HC(\mathcal{O}_{2})$. Consequently, $HC(\mathcal{O}_{2})$ could have two ergodic measures of maximal entropy.
2.3. Partially hyperbolic systems
@@ 2.3.1. Existence of equilibrium states
$P_{metric}(f,\phi):{\mathcal{M}}_{f}\to\mathbb{R},\,\,\mu\mapsto P_{\mu}(f,\phi).$
@@ 2.4. Unstable entropy
In this section, we recall the notions of unstable $h^{u}(f)$ and stable $h^{s}(f)$ topological entropy for a partially hyperbolic diffeomorphism $f$, see [53, 17, 49, 22, 23].
@@ 2.4.1. Unstable metric entropy
Consider a compact metric space $M$ with the Borel $\sigma$-algebra ${{\mathcal{B}}}$ and let $\mu$ represent a Borel probability measure defined on this space. Let ${{\mathcal{P}}}=\{P_{i}\}_{i\in I}$ be a partition of $M$ into measurable subsets, i.e., $M=\bigcup_{P_{i}\in\mathcal{P}}P_{i}$ such that $P_{i}\in{{\mathcal{B}}}$, $\mu(P_{i}\cap P_{j})=0$, and $\mu\left(\bigcup_{P_{i}\in{{\mathcal{P}}}}P_{i}\right)=\mu(M)=1$. We denote by $\sigma({{\mathcal{P}}})$ the smallest $\sigma$-algebra containing all $P_{i}\in{{\mathcal{P}}}$.
@@
We say that a partition ${{\mathcal{P}}}$ is measurable (or countably generated) with respect to $\mu$ if there exists a family of measurable sets $\{A_{i}\}_{i\in\mathbb{N}}$ and a measurable set $F$ of full measure such that if $B\in{{\mathcal{P}}}$, then there exists a sequence $\{B_{i}\}$, where $B_{i}\in\{A_{i},A_{i}^{c}\}$ such that $B\cap F=\bigcap_{i}(B_{i}\cap F)$.
Consider $\pi:M\rightarrow M/\mathcal{P}$ as the canonical projection, mapping each point in $M$ to the corresponding element in the quotient space formed by the atoms of the partition. This naturally gives rise to a measurable structure on $M/\mathcal{P}$ via the measure $\widetilde{\mu}=\pi^{*}\mu$ defining it on the associated pushed $\sigma$-algebra. Additionally, for any point $x$ in $M$, $P(x)$ represents the element of the partition to which $x$ belongs.
@@
[42, Rokhlin’s Disintegration] Let $P$ be a measurable partition of a compact metric space $M$, and $\mu$ a Borel probability measure. Then there exists a unique disintegration by conditional probability measures for $\mu$.
@@
$h_{\mu}(\mathcal{P},f)=\lim_{n\to\infty}\frac{1}{n}H_{\mu}(\mathcal{P}_{n})=\lim_{n\to\infty}\left(H_{\mu}(\mathcal{P}_{n})-H_{\mu}(\mathcal{P}_{n-1})\right),$
where $\mathcal{P}_{n}=\mathcal{P}\vee f^{-1}(\mathcal{P})\vee\ldots\vee f^{-n+1}(\mathcal{P})$ is a refinement of $\mathcal{P}$ by the action of $f$.
Now let ${{\mathcal{F}}}$ be an $f$-invariant foliation, that is, $f({{\mathcal{F}}}(x))={{\mathcal{F}}}(f(x))$, for every $x\in M$, where $\mathcal{F}(x)$ denotes the leaf at $x$.
We say that ${{\mathcal{F}}}$ is uniformly expanding if there exists $\lambda>1$ such that $\forall x\in M$, $\|Df^{-1}_{|T_{x}(F(x))}\|<\lambda^{-1}$. A measurable partition $\xi$ is called increasing and subordinated to ${{\mathcal{F}}}$ if it satisfies the following properties:
@@
Finding a measurable and increasing partition that aligns with an invariant lamination is typically a complex challenge. However, when dealing with a uniformly expanding foliation that remains invariant under a diffeomorphism, such a partition is always available [53, 22, 23]. Additionally, for expanding foliations ${{\mathcal{F}}}$ it holds
@@
It can be proved that the above definition is independent of $\xi$ [17, Lemma 2.8].
@@
$h^{s}_{\mu}(f):=h_{\mu}(f^{-1},{{\mathcal{F}}}^{s}),$
where ${{\mathcal{F}}}^{s}$ denotes the stable foliation given by the stable manifolds.